"The magnitude of the force between two charges of 1.5µc each that 5cm apart."


1
Expert's answer
2022-06-01T12:49:13-0400
"F=k\frac{q_1q_2}{r^2}=9*10^9*\frac{(1.5*10^{-6})^2}{0.05^2}=8.1\:\rm N"